A whiskey 0.789 gml, and O997 g℡L, 4, wtains 40 0% cthanol. Cillio I. by volume. The desitles of ethanol and water are 40.0 respectively, Calculate the ethanol concentrations in whiskey. 110 pts (a) mass percent of ethanol (b) mole fraction of ethanol (MWe 46.07 g/mol and MWe 18.016 g/mol) (e) molarity of ethanol d) molality of ethanol
Answer #1

(a) 40 % ethanol by volume meaning 100 ml solution contain 40 ml ethanol and 60 ml water

mass = density X volume

mass of ethanol in 100 ml solution = 0.789 X 40 = 31.56 gm

mass of water in 100 ml solution = 0.997 X 60 = 59.82 gm

total mass of solution = 31.56 + 59.82 = 91.38 gm

mass % of solue = mass of solute X 100 / mass of solution

mass % of ethanol = 31.56 X 100 / 91.38 = 34.54 %

mass % of ethanol = 34.54 %

(b) molar mass of ethanol = 46.07 g/mol then 31.56 gm of ethanol = 31.56/46.07 = 0.685 mole

molar mass of water = 18.016 g/mole then 59.82 gm of water = 59.82 / 18.016 = 3.32 mole

total mole = 0.685 + 3.32 = 4.005 mole

mole fraction of solute = no, of mole of of solute / total mole

mole fraction of ethanol = 0.685/4.005 = 0.171

mole fraction of ethanol = 0.171

(c) molariy = no. of mole / volume of soltion in liter

no.of mole of ethanol = 0.685 mole

volume of solution = 100 ml = 0.1 liter

molarity of ethanol = 0.685 / 0.1 = 6.85 M

molarity of ethanol = 6.85 M

(d) molality = no. of mole of solute / mass of solution in Kg

no.of mole of ethanol = 0.685 mole

mass of solution = 91.38 gm = 0.09138 kg

molality of etthanol = 0.685/91.38 = 7.50 m

molality of etthanol = 7.50 m